021 (part 1 of 2) 10.0 points A student sits on a rotating stool holding two 4 kg objects. When his arms are extended horizontally, the objects are 0.7 m from the axis of rotation, and he rotates with angular speed of 0.69 rad/sec. The moment of iner- tia of the student plus the stool is 4 kg and is a ssumed to be constant. The student then pulls the objects horizontally to a radius 0.29 m from the rotation axis. Calculate the final angular speed of the student Answer in units of rad/s. 022 (part 2 of 2) 10.0 points Caleulate the change in kinetic energy of the system. Answer in units of J.
Answer #1

Part 1 of 2 .

Moment of inertia of Student with Stool ; Is = 4 kg-m2

When there are 2 objects of 4 kg each at radius 0.7m, it will add moment of inertia of the system; so, total moment of inertia will be

I1 = Is + 2* 4 * (0.7)2 = 4 + 8* 0.49 = 7.92 kg-m2

Angular speed ω1 = 0.69 rad/s

When arms are pulled and objects are at 0.29 m distance,  total moment of inertia will be

I2 = Is + 2* 4 * (0.29)2 = 4 + 8* 0.0841 = 4.6728 kg-m2

As the angular momentum is conserved , we will have

I1ω1 = I2ω2 ; where ω2 is Angular speed in second case

ω2 = I1ω1  /I2 = 7.92 * 0.69 / 4.6728 = 1.17 rad/s

Part 2 of 2

Change in Kinetic energy will be given as

(1/2) I2ω22 - (1/2) I1ω12

= (1/2) * 4.6728 * 1.172 - (1/2) * 7.92 * 0.692

= 1.31 J